OpenStudy (anonymous):

what is the product of 2.4 times 10 cubes and 1.5 times 10 cubed

OpenStudy (anonymous):

(2.4)*(10^3)= 2.4*1000=2400, not u try the next one.

OpenStudy (anonymous):

If u have a calculator, it would be easier to compute.